OpenStudy (anonymous):

One more rectangular-shaped piece of metal siding needs to be cut to cover the exterior of a pole barn. The area of the piece is 30 ft². The length is 1 less than 3 times the width. How wide should the metal piece be? Round to the nearest hundredth of a foot. (1 point) 3.33 ft 4.3 ft 1 ft 30 ft

OpenStudy (anonymous):

LW = (3W-1) W = 30 3W^2 - W - 30=0 W=(1+/- sqrt(1-4*3*-30))/6 = (1+/-sqrt(1+12*30))/6 = (1+/- sqrt(361))/6 = (1+19)/6 = 20/6 ~ = 3.33